# The Complexity of HyperQPTL

### v2

@@ 5 Conclusion
We have settled the exact complexity of the most important verification problems for $\mathrm{Hyper\-QPTL}$ and $\mathrm{Hyper\-QPTL^{+}}$. For $\mathrm{Hyper\-QPTL}$, we proved that satisfiability is $\Sigma^{2}_{1}$-complete while for $\mathrm{Hyper\-QPTL^{+}}$, we proved that satisfiability, finite-state satisfiability, and model-checking are equivalent to truth in third-order arithmetic. The latter results were obtained by showing that $\mathrm{Hyper\-QPTL^{+}}$ and second-order $\mathrm{Hyper\-LTL}$ have the same expressiveness.
Acknowledgments
Supported by DIREC – Digital Research Centre Denmark.

### v3

@@ 5 Conclusion
We settled the exact complexity of the most important verification problems for $\mathrm{Hyper\-QPTL}$ and $\mathrm{Hyper\-QPTL^{+}}$. For $\mathrm{Hyper\-QPTL}$, we proved that satisfiability is $\Sigma^{2}_{1}$-complete while for $\mathrm{Hyper\-QPTL^{+}}$, we proved that satisfiability, finite-state satisfiability, and model-checking are equivalent to truth in third-order arithmetic. The latter results were obtained by showing that $\mathrm{Hyper\-QPTL^{+}}$ and second-order $\mathrm{Hyper\-LTL}$ have the same expressiveness.
Acknowledgments. Gaëtan Regaud was supported by the European Union. Martin Zimmermann was supported by DIREC – Digital Research Centre Denmark. We thank the reviewers for their detailed comments, which improved the article considerably.